OpenStudy (anonymous):

what is the equation in y=mx+b form, where the line passes through (-2,1) and is parallel to 4x-3y-7=0 ?

OpenStudy (anonymous):

If its parallel to 4x-3y-7=0 then it has the same slope as that line.

OpenStudy (anonymous):

Solve 4x-3y-7=0 for Y

OpenStudy (anonymous):

4x-3y = 7

OpenStudy (anonymous):

-3Y = -4X + 7

OpenStudy (anonymous):

okay, got it, thank you!

OpenStudy (anonymous):

What did you get for an answer?